02 1 2. X= 1 1 -2 X 2 2 -1 1 2 0 3. X = X+ 1 e tant 2

number 3 use variation of parameters for homogeneous linear systems
Answer #1

Page No.: ار 0 4 مد X =) x= AX 4 4 - = ه و وسه 1 To find eixen Value cha (A) 4 4 - A A-de 3 1 و \ وس - -d +46 (3-به-2ص) : (هT ४५ - x 1 HA नि nz 4. ----- 31 1-1- ; (A-da I)....o 2-| O 2 10 C -2 2-2 00 LE . 0 . a) ४५ - 20 + । Hong Vig thea Om/ HAK-4 g1 2 X + les seves tout AX + git Ao fi :] gitla | et tont - A-11 The S= }*-**+2=0 i = 1+i die Hi . do = 1+1 for 1,1-1 IA-2,71V=et [ Cost - Sint) i sine ( ) 2 Sint een tiet/ 2 Cost Cost - Sint Mi1t) = V+ iv 2 lost al sind 2 sint Cost el - sint GenerelSint -Sint tant Now Sint - In tontt somnade xija Cust tent -Sint tant Sint s sint at s-sint tant de {xt e Elle 4) и -Sint tan